OpenStudy (anonymous):

Will medal and fan! Use the point slope form, write an equation for the line perpendicular to the line y = -2(x - 5) containing the point (3,4).

OpenStudy (anonymous):

Point slope form: \[y-y _{1}=m(x-x _{1})\]

OpenStudy (jdoe0001):

any ideas on what the slope of \(\bf y = -2(x - 5)\) is?

OpenStudy (anonymous):

-2 is the slope right?

OpenStudy (anonymous):

@jdoe0001

Nnesha (nnesha):

yep that's right :

OpenStudy (anonymous):

So what's the next step?

Nnesha (nnesha):

perpendicular slopes are negative reciprocal if slope of first equation is a/b then slope of perpendicular line would be negative b/a

OpenStudy (anonymous):

So it would be -1/2?

Nnesha (nnesha):

hmm no.

OpenStudy (jdoe0001):

slope is -2/ a perpendicular line to that one, as Nnesha said, would have a NEGATIVE RECIPROCAL SLOPE, or \(\bf slope=-2\qquad negative\implies +2\qquad reciprocal\implies +\cfrac{1}{2}\)

Nnesha (nnesha):

^

OpenStudy (anonymous):

Oh positive 1/2! whoops

OpenStudy (anonymous):

So once I find that then what?

OpenStudy (jdoe0001):

so, you're really looking for the equation of a line that has a slope of 1/2 and passes through (3,4) \(\bf \begin{array}{lllll} &x_1&y_1\\ % (a,b) &({\color{red}{ 3}}\quad ,&{\color{blue}{ 4}})\quad \end{array} \\\quad \\ % slope = m slope = {\color{green}{ m}}=\cfrac{1}{2} \\ \quad \\ % point-slope intercept y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{plug in the values and solve for "y"}\\ \qquad \uparrow\\ \textit{point-slope form}\)

OpenStudy (anonymous):

So that would equal y - 4 = 1/2x - 3/2

OpenStudy (anonymous):

y = 1/2x +5/2 would be the answer right?

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

That's the answer? Sorry I wasn't sure which reply you were saying yes to.

OpenStudy (anonymous):

@jdoe0001

OpenStudy (anonymous):

Okay thanks!

OpenStudy (jdoe0001):

even more typos

OpenStudy (jdoe0001):

\(\bf y-4=\cfrac{1}{2}(x-3)\implies y=\cfrac{1}{2}x-\cfrac{3}{2}+4\implies y=\cfrac{1}{2}x-\cfrac{3}{2}+\cfrac{8}{2} \\ \quad \\y=\cfrac{1}{2}x+\cfrac{5}{2}\)

OpenStudy (anonymous):

Lol it's okay. I got it! :)

OpenStudy (jdoe0001):

:)

Nnesha (nnesha):

o^_^o